Status Report

NOAA SATOPS Morning Report: Monday, May 21, 2012

By SpaceRef Editor
May 21, 2012
Filed under ,

SATOPS Morning Report: Monday, May 21, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 142

GOES
GOES-12 DOY142 intermittent Imager instrument cycle slips from 03:22z through 12:41z.

GOES-12 DOY 141 intermittent Imager instrument cycle slips from 09:16z through 11:42z.

All other GOES operations were nominal over the past 72 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
MODIS data outage from May 20,2012 20:00z to May 20, 2012 21:00z

Metop data between sensing time DOY140 12:08 and 14:44, orbits #28951 and #28952 were delayed due to other ADA ground station priories.

Questions or comments to [email protected]

SpaceRef staff editor.